OpenStudy (anonymous):

Find lim x as approaches 0 sin(x)/(2x^2-x). Use only limits properties or sandwich theorem,since those are the only rules I have learnt so far;)

OpenStudy (anonymous):

\[\LARGE \lim_{x\to 0}\frac{\sin (x)}{x}=?\] this rule is equal to what? :)

OpenStudy (zarkon):

\[\frac{\sin(x)}{2x^2-x}=\frac{\sin(x)}{x(2x-1)}=\frac{\sin(x)}{x}\frac{1}{2x-1}\]

OpenStudy (anonymous):

I see

OpenStudy (anonymous):

Thanks guys